# Recurrent Aligned Network for Generalized Pedestrian Trajectory Prediction

### v1

@@ 2 $S_{i,t}\leftarrow h_{i,t}^{n}$triangleright Equation 6
4       $\mathcal{L}_{rec}(t)\leftarrow[S_{1,t},S_{2,t}]$triangleright Equation 7
5
@@ 8 Backward Loss $\mathcal{L}_{rec}$
where $h_{i,t}^{n}$ is the hidden state of the trajectory’s time-state representation $X_{i,t}^{n}$. $r_{i,t}^{n}$, $z_{i,t}^{n}$, and $g_{i,t}^{n}$ are the reset, update, and new gates, respectively. $W$ and $b$ denote learning parameters. $\sigma$ is the activation function. $\odot$ is the element-wise multiplication. The initial hidden state $h_{i,0}$ is produced by a linear transformation of the feature embedding $\mathbf{F}_{i,t}^{b}$. Hence, the feature space of each domain at different timestamps can be represented as:
$\begin{split}S_{i,t}=\sum_{n=1}^{N_{i}}h_{i,t}^{n},\end{split}$ (6)
where $S_{i,t}$ is the feature space of all trajectories in domain $i$ at timestamp $t$. Given $S_{i,t}$, we can minimize the discrepancy between the two source domains using the proposed recurrent alignment strategy and the recurrent alignment losses outlined in Algorithm 1. Then, our network is capable of modeling generalized trajectory features, which can be regarded as domain-invariant. Recurrent Alignment Loss. The alignment loss is to minimize the discrepancy between source domains. To support the proposed recurrent alignment strategy, we design a recurrent alignment loss to minimize the discrepancy between the two source domains at both time-state and time-sequence levels as follows:
@@
where $\mathcal{L}_{rec}$ is the recurrent alignment loss, which calculates the total discrepancy between the two source domains at each timestamp through the recurrent alignment strategy. $||S_{1,t},S_{2,t}||_{\mathcal{H}}$ is the alignment loss at timestamp $t$, indicating the discrepancy measurement between two source domains. There are multiple choices for measuring such discrepancies, such as $L_{2}$ distance, MMD, and CORAL. We explore six measurement functions in Section 4 and choose $L_{2}$ distance loss in this work.
3.4 Prediction and Loss Function
Pedestrian trajectory prediction aims to predict future trajectories that closely match the ground truth. Through the pre-aligned representation module and the recurrent alignment module, we can obtain the generalized trajectory features. Then, we need to decode these features into predictions as follows:
@@ 3.4 Prediction and Loss Function
where $(\mathbf{\hat{Y}}_{i}^{n})^{K}$ is $K$ predictions of the $n^{th}$ pedestrian in domain $i$ following previous best-of-$K$ metrics. $f_{d}$ is a mixture of expert decoders, implemented by $K$ MLP layers. We use the $L_{2}$ distance loss as the prediction loss as follows:
@@
where $\mathcal{L}_{pre}$ is the prediction loss of all trajectories in all source domains. $(\mathbf{\hat{Y}}_{i}^{n})^{min}$ denotes the prediction with the minimum error in $(\mathbf{\hat{Y}}_{i}^{n})^{k}$. The overall network is trained end-to-end by minimizing the total loss function $\mathcal{L}_{RAN}$ as follows:
@@
Figure 3: The inference phase of the proposed RAN framework.
3.5 Inference
@@
Because the two RNN layers share weights and the two stepwise attention layers share weights during training, the inference phase is straightforward, as illustrated in Figure 3. Given an observed trajectory from the target domain, we initially model its stepwise interaction using the stepwise attention layer. Subsequently, we combine this attention information with the trajectory embedding and input it into the well-aligned RNN layer to extract generalized trajectory features. Finally, we employ a multimodal trajectory decoder to decode the generalized trajectory features into the predicted trajectories.
4 Experiments and Discussions
4.1 Evaluation Datasets.
We evaluate our proposed method on three widely used benchmark datasets, each situated in different scenarios and representing different domains. ETH-UCY. The ETH-UCY dataset  [17, 27] is captured in the scenes of outdoor pedestrian paths. It contains trajectories of 1,536 pedestrians in five subsets, i.e., ETH, HOTEL, UNIV, ZARA1, and ZARA2. The trajectories within the ETH-UCY dataset are documented in world coordinates, and thus we report the results in meters as previous methods [40, 1]. SDD. The Stanford Drone Datase (SDD)  [29] is captured in the scenes of university campus. It is a well established benchmark for human trajectory prediction in bird’s eye view, which consists of 20 subsets captured use a drone. Trajectories in SDD are documented in camera coordinates, and hence, we report the results in pixels as previous approaches [40, 1]. NBA. The SportVU NBA movement (NBA) dataset  [43] is captured in the scenes of basketball courts. It is a more challenging benchmark that contains rich social interactions. Due to the large size of the original dataset, we randomly extracted 257,230 trajectories in the “Rebound” subset following the previous method [40]. Trajectories in NBA dataset are documented in world coordinates, and hence, we report the results in meters as the previous approach [40].
4.2 Experimental Setting
Evaluation Protocol. We treat the three datasets, i.e., ETH-UCY, SDD and NBA datasets, in different scenarios as different domains. For generalized pedestrian trajectory prediction, the model can not reach the target domain data in the training phase. Specifically, taking source domains A and B and target domains C as examples, we split A, B and C into train sets and test sets, respectively. All train sets and the test set do not overlap. We train and validate the model on train and test sets of A and B, respectively. Then, we test the model on the test set of C. More details are given in the supplemental material. Evaluation Metrics. Following previous methods [40, 41], we employ two commonly used metrics, i.e., Average Displacement Error (ADE) and Final Displacement Error (FDE), as our evaluation metrics. ADE means the average $L_{2}$ distance between all points of the prediction and ground truth. FDE means the $L_{2}$ distance between the destination points of the prediction and ground truth. Implementation Details. Similar with the previous methods [41, 40, 26], we calculate the best-of-20 performance. We observe 8 frames for each trajectory and predict the subsequent 12 frames. The two MLP encoders are all 2-layer linear transformations with the sigmoid activation function, and the hidden sizes are set to 64 and 128. The dimensions of the learnable parameters $W_{i,t}^{q}$, $W_{i,t}^{k}$, and $W_{i,t}^{v}$ in attention are all set to 256. We employ GRUs as the RNN structure. The activation function $\sigma$ is sigmoid. The coefficients $\lambda_{1}$ and $\lambda_{2}$ in the loss function are set to 1. The Adam optimizer is used to train our model by 300 epochs with a learning rate of 0.001 and batch size of 512, decaying by 0.5 with an interval of 50, on one GTX-3090 GPU. We will release the code later for details.
@@ 4.2 Experimental Setting
Table 1: Comparison for generalization results (ADE/FDE) on ETH-UCY. Model is trained on SDD and NBA. $\ast$ means the results are reproduced using the official released code. Bold indicates the best performance. The Lower the better.
@@
Target Domain Social-STG ${}^{\ast}$ [26] ${}^{\ast}$ [31] ${}^{\ast}$ [23] GP- ${}^{\ast}$ [2] Social- ${}^{\ast}$ [40] Graph- ${}^{\ast}$ [1] Ours
CVPR2020 CVPR2021 AAAI2021 ECCV2022 ECCV2022 AAAI2023
@@
4.3 Quantitative Analysis
Tables 1-3 show the performance compared to six state-of-the-art methods in three widely used datasets, which are regarded as three different domains. The model is trained on two selected domains and tested on the remaining one to evaluate cross-domain generalization. Consistent with established protocols, we express ETH-UCY and NBA results in meters and SDD results in pixels. To eliminate the impact of different units of measurement on the model, we unified them as meters during training. We reproduced previous state-of-the-art methods under the generalization setting through their official released codes without post-processing to ensure fair comparisons. ETH-UCY. Table 1 shows the generalization performance on the ETH-UCY dataset. It shows that our method achieves the best performance in both ADE and FDE. Specifically, our method improves the average ADE/FDE from $0.36/0.60$ to $0.23/0.42$ compared to the previous best performance method, Graph-TERN [1]. Besides, we also outperform Graph-TERN in all five subsets, which demonstrates the superior generalization capability of our approach. SDD. Table 2 shows the generalization performance on the SDD dataset. It shows that our method achieves the best performance in both ADE and FDE compared with the other methods. Specifically, our method improves the average ADE/FDE from $10.97/19.95$ to $14.76/25.47$ compared to the previous best performance method, Graph-TERN [1]. NBA. Table 3 shows the generalization performance on the NBA dataset. It shows that our method achieves the best performance in both ADE and FDE. Specifically, our method improves the average ADE/FDE from $1.28/2.64$ to $2.11/4.45$ compared to the previous best performance method, SocialVAE [40]. Notably, NBA dataset is captured in the basketball court, which presents denser interactions and a larger domain gap compared to the other datasets. Our method achieve more performance improvement facing such larger domain gap, i.e. $44.8\%/45.3\%$ (NBA), $36.1\%/30\%$ (ETH-UCY) and $25.6\%/21.6\%$ (SDD), compared with Graph-TERN [1]. These results underscore our method’s exceptional ability to generalize across varied and complex domains.
Table 2: Comparison for generalization results (ADE/FDE) on SDD. Model is trained on ETH-UCY and NBA. $\ast$ means the results are reproduced using the official released code. Bold indicates the best performance. The Lower the better.
Target Domain Social-STG ${}^{\ast}$ [26] ${}^{\ast}$ [31] ${}^{\ast}$ [23] GP- ${}^{\ast}$ [2] Social- ${}^{\ast}$ [40] Graph- ${}^{\ast}$ [1] Ours
@@ 4.3 Quantitative Analysis
Table 3: Comparison for generalization results (ADE/FDE) on NBA. Model is trained on ETH-UCY and SDD. $\ast$ means the results are reproduced using the official released code. Bold indicates the best performance. The Lower the better.
Target Domain Social-STG ${}^{\ast}$ [26] ${}^{\ast}$ [31] ${}^{\ast}$ [23] GP- ${}^{\ast}$ [2] Social- ${}^{\ast}$ [40] Graph- ${}^{\ast}$ [1] Ours
@@
Table 4: Ablation study of each components of our method in ADE/FDE in meters on ETH-UCY. The lower the better.
@@
Table 5: Ablation study of recurrent alignment loss on ETH-UCY dataset.
@@
4.4 Ablation Study
In this section, we first study the contribution of the important modules of RAN. Then we evaluate the impact of different alignment approaches. Finally, we study the impact of different alignment units and coefficients of our loss function. Contribution of Each Component. We conduct two variants to evaluate the components of our RAN as shown in Table 4. PAR and RA denote the pre-aligned representation module and the recurrent alignment module, respectively. PAR is replaced by only embedding the target trajectory for recurrent alignment at each timestamp. RA is removed to evaluate our recurrent aligned strategy. The experimental results show that each component is effective in improving the generalization capability of our method. Analysis of Different Alignment Approaches. As shown in Figure 4, we replace our recurrent alignment strategy in RAN with time-state and time-sequence alignment on the ETH-UCY dataset. “State” denotes time-state alignment, implemented by aligning the feature space at each step at attention layers. “Sequence” denotes time-sequence alignment, implemented by only aligning the feature space at the last step at RNN layers. “Recurrent” denotes our recurrent alignment strategy, which considers both time-state and time-sequence alignment. Experiment results show that our recurrent alignment strategy achieves the best performance, which indicates the superiority of our method in domain alignment for generalized pedestrian trajectory prediction. Analysis of Recurrent Alignment Loss. As shown in Table 5, we conduct experiments to evaluate the impact of using different alignment losses as the alignment unit in $\mathcal{L}_{rec}$. Specifically, MMD, CORAL, KLD, JS, and COS denote using MMD loss, CORAL loss, Kullback-Leibler divergence loss, Jensen–Shannon divergence loss, and cosine similarity as the alignment unit in $\mathcal{L}_{rec}$, respectively. As mentioned in Section 3, we use the $L_{2}$ distance as the alignment unit in $\mathcal{L}_{rec}$. The evaluation results suggest that the $L_{2}$ distance is a more suitable domain distance measurement metric for pedestrian trajectory prediction tasks. One possible explanation is that high-dimensional feature representations may still preserve spatial-level information, i.e., the spatial distance, in pedestrian trajectory prediction tasks. Analysis of Loss Coefficient. As shown in Figure 5, we evaluate different loss coefficients to validate the appropriate proportion of the alignment module in our framework. The proportion of the alignment module is represented by the figure’s loss coefficient ratio $C=\lambda_{1}/\lambda_{2}$. The evaluation results indicate that an excessive proportion of alignment modules sacrifices the model’s prediction accuracy, resulting in a performance decline. Conversely, a deficient proportion of alignment modules diminishes the model’s generalization capacity, leading to a performance decrease. Based on the experimental results, we choose $C=1$ in this work, which achieves the best performance.

### v2

@@ 3 $S_{i,t}\leftarrow h_{i,t}^{n}$Equation 6}
5       $\mathcal{L}_{rec}(t)\leftarrow[S_{1,t},S_{2,t}]$Equation 11}
6
@@ 9 Backward Loss $\mathcal{L}_{rec}$
Recurrent Alignment Loss. The alignment loss is to minimize the discrepancy between the source domains. To support the proposed recurrent alignment strategy, we design a recurrent alignment loss to minimize the discrepancy between the two source domains at both time-state and time-sequence levels as follows:
@@
where $\mathcal{L}_{rec}$ is the recurrent alignment loss, which calculates the total discrepancy between the two source domains at each time step through the recurrent alignment strategy. $||S_{1,t},S_{2,t}||_{\mathcal{H}}$ is the alignment loss at time step $t$, indicating the discrepancy measurement between two source domains. There are multiple choices for measuring such discrepancies, such as $L_{2}$ distance, MMD, and CORAL. We explore six measurement functions in Section IV and choose $L_{2}$ distance loss for the generalized pedestrian trajectory prediction in this work.
Figure 3: The inference phase of the proposed RAN framework. The input data are from target domain scenarios, and there is no requirement for data from at least two domains during the inference phase.
III-D Prediction and Loss Function
Pedestrian trajectory prediction aims to predict future trajectories that closely match the ground truth. Due to the multiple possibilities in the future, it is reasonable to have multiple potential future trajectories for a single observed trajectory. Following the previous best-of-K protocol, we use the mixture of expert (MoE) decoders to generate the corresponding $K$ possible predictive trajectories and choose the prediction that most closely matches the ground truth to calculate performance. Through the pre-aligned representation module and the recurrent alignment module, we can obtain the generalized trajectory features. Then, we need to decode these features into predictions as follows:
@@ III-D Prediction and Loss Function
where $(\mathbf{\hat{Y}}_{i}^{n})^{K}$ is $K$ predictions of the $n^{th}$ pedestrian in domain $i$ following previous best-of-$K$ metrics. $f_{d}$ is the mixture of expert decoders implemented by $K$ MLP layers. We use the $L_{2}$ distance loss as the prediction loss as follows:
@@
where $\mathcal{L}_{pre}$ is the prediction loss of all trajectories in all source domains. $(\mathbf{\hat{Y}}_{i}^{n})^{min}$ denotes the prediction with the minimum error in $(\mathbf{\hat{Y}}_{i}^{n})^{k}$.
The overall network is trained end-to-end by minimizing the total loss function $\mathcal{L}_{RAN}$ as follows:
@@
III-E Inference
@@
Because the two RNN layers share weights and the two stepwise attention layers share weights during training, the inference phase is straightforward, as illustrated in Figure 3. Given an observed trajectory from the target domain scenarios, we initially model its stepwise interaction at each time state using the stepwise attention layer. Subsequently, we combine this attention information with the trajectory embedding and input it into the well-aligned RNN layer to extract generalized trajectory features. Finally, we employ a multimodal trajectory decoder to decode the generalized trajectory features into the multimodal future predictions.
III-F Multiple Source Domains Discussion
Theoretically, our method can handle an unlimited number of source domains in training by adapting the recurrent alignment loss. Specifically, when the number of source domains exceeds two, the recurrent alignment loss $\mathcal{L}_{recm}$ is calculated as follows:
$\begin{split}\mathcal{L}_{recm}&=\frac{1}{m^{2}}\sum_{t=1}^{T_{obs}}\sum_{p,q=1}^{m}||S_{p,t},S_{q,t}||_{\mathcal{H}},\\ \end{split}$ (11)
where $||S_{p,t},S_{q,t}||_{\mathcal{H}}$ is the alignment loss between two source domains $p$ and $q$ at time step $t$. $m$ is the number of source domains. However, as the number of source domains increases, the computation complexity of the recurrent alignment loss also increases. Specifically, during training, if there are $m$ source domains, and each source domain has a trajectory observation time of $T_{obs}$, then the computational complexity of the recurrent alignment loss $\mathcal{L}_{recm}$ is $O(T_{obs}\cdot m^{2})$.
IV Experiments and Discussions
@@
TABLE II: Comparison for generalization results (ADE/FDE) on ETH-UCY. The model is trained in SDD and NBA. $\ast$ means the results are reproduced using the officially released code. Bold indicates the best performance. The lower the better.
@@
Target Domain Social-STG CNN [61] SGCN [22] TPNMS [62] GP- Graph [24] Social- VAE [1] Graph- TERN [63] Ours
@@
TABLE III: Comparison for generalization results (ADE/FDE) on SDD. The model is trained on ETH-UCY and NBA. $\ast$ means the results are reproduced using the official released code. Bold indicates the best performance. The Lower the better.
Target Domain Social-STG CNN [61] SGCN [22] TPNMS [62] GP- Graph [24] Social- VAE [1] Graph- TERN [63] Ours
@@
TABLE IV: Dataset split for train/test/whole sets for three different datasets. Trajectories in the train set and test set do not overlap.
Dataset Number of Trajectories
Training Set Test Set Whole Set
ETH-UCY 34,161 3,109 37,270
SDD 8,494 2,829 11,323
NBA 205,970 51,260 257,230
In this section, we evaluate the pedestrian trajectory prediction performance of our proposed RAN and compare it with existing state-of-the-art methods on four standard benchmark datasets. Additionally, we conduct detailed ablation studies to explore the contribution of each component.
IV-A Evaluation Datasets.
We evaluate our proposed method on three widely used benchmark datasets, each situated in different scenarios and representing different domains.
ETH-UCY. The ETH-UCY dataset  [64, 65] is captured in the scenes of outdoor pedestrian paths. It contains trajectories of 1,536 pedestrians in four different scenes, i.e., ETH, HOTEL, UNIV, ZARA. It contains five sub-datasets captured from downfacing surveillance cameras annotated at 2.5Hz. To ensure a fair evaluation, we follow the previous leave-one-out protocol widely adopted in previous studies [1, 63], where it is trained on four subsets and tested on the remaining one. The trajectories within the ETH-UCY dataset are documented in world coordinates, and thus we report the results in meters as previous methods [1, 63].
SDD. The Stanford Drone Datase (SDD)  [66] is captured in the scenes of the university campus. It is a well-established benchmark for human trajectory prediction in bird’s eye view, which consists of 20 subsets captured using a drone. It collects multi-agent trajectories of various types, including pedestrians, bicyclists, skateboarders, cars, buses, and golf carts, in a university campus setting. The dataset is notable for its vast size, containing over 11,000 individual pedestrian trajectories, which result in more than 185,000 interactions between pedestrians and over 40,000 interactions between pedestrians and their surroundings. To ensure a fair and consistent evaluation of our method, we adhere to the standard training and testing splits used in previous studies [1]. Trajectories in SDD are documented in camera coordinates, and hence, we report the results in pixels as previous approaches [1, 63].
NBA. The SportVU NBA movement (NBA) dataset  [67] is captured in the scenes of basketball courts. It is a more challenging benchmark that contains rich social interactions. Due to the large size of the original dataset, we randomly extracted 257,230 trajectories in the “Rebound” subset following the previous method [1]. The extracted trajectories capture a rich set of agent-agent interactions and highly non-linear motions. It is important to note that the overall frequency and the adversarial and cooperative nature of the interactions in this dataset are significantly different from those in ETH/UCY and SDD [1]. Trajectories in the NBA dataset are documented in world coordinates, and hence, we report the results in meters as the previous approach [1].
TABLE V: Comparison for generalization results (ADE/FDE) on NBA. The model is trained on ETH-UCY and SDD. $\ast$ means the results are reproduced using the official released code. Bold indicates the best performance. The Lower the better.
Target Domain Social-STG CNN [61] SGCN [22] TPNMS [62] GP- Graph [24] Social- VAE [1] Graph- TERN [63] Ours
@@
TABLE VI: Ablation study of each component of our method in ADE/FDE in meters on ETH-UCY. The model is trained in SDD and NBA. The lower the better.
@@
IV-B Experimental Setting
Evaluation Protocol. We treat the three datasets, i.e., ETH-UCY, SDD and NBA datasets, in different scenarios as different domains. For generalized pedestrian trajectory prediction, the model can not reach the target domain data in the training phase. Specifically, taking source domains A and B and target domains C as examples, we split A, B and C into train sets and test sets, respectively. All train sets and the test set do not overlap. We train and validate the model on train and test sets of A and B, respectively. Then, we test the model on the test set of C. More details can be found in Tabble IV.
Evaluation Metrics. Following previous methods [1, 14], we employ two commonly used metrics, i.e., Average Displacement Error (ADE) and Final Displacement Error (FDE), as our evaluation metrics. ADE means the average $L_{2}$ distance between all points of the prediction and ground truth. FDE means the $L_{2}$ distance between the destination points of the prediction and ground truth.
Implementation Details. Similar with the previous methods [14, 1, 61], we calculate the best-of-20 performance. We observe 8 frames for each trajectory and predict the subsequent 12 frames. The two MLP encoders are all 2-layer linear transformations with the sigmoid activation function, and the hidden sizes are set to 64 and 128. The dimensions of the learnable parameters $W_{i,t}^{q}$, $W_{i,t}^{k}$, and $W_{i,t}^{v}$ in attention are all set to 256. We employ GRUs as the RNN structure. The activation function $\sigma$ is sigmoid. The coefficients $\lambda_{1}$ and $\lambda_{2}$ in the loss function are set to 1. The Adam optimizer is used to train our model by 300 epochs with a learning rate of 0.001 and batch size of 512, decaying by 0.5 with an interval of 50, on one GTX-3090 GPU.
TABLE VII: Comparison for performance (ADE/FDE) on ETH-UCY. Results are trained on different source domains and tested on the target domain ETH-UCY. Bold indicates the best performance. The lower the better.
Social-STG CNN [61] SGCN [22] TPNMS [62] GP- Graph [24] Social- VAE [1] Graph- TERN [63] Ours
Case 1 Source: ETH-UCY Target: ETH-UCY 0.44/0.75 0.37/0.65 0.38/0.73 0.23/0.39 0.24/0.42 0.24/0.38 0.21/0.36
Case 2 Source: ETH-UCY & SDD Target: ETH-UCY 0.44/0.75 0.35/0.65 0.38/0.73 0.23/0.39 0.23/0.41 0.23/0.38 0.20/0.36
Case 3 Source: NBA & SDD Target: ETH-UCY 0.64/0.97 0.57/0.88 0.63/0.94 0.37/0.62 0.49/0.72 0.36/0.60 0.23/0.42
TABLE VIII: Ablation study of recurrent alignment loss on ETH-UCY dataset. The model is trained in SDD and NBA. The lower the better.
@@ IV-B Experimental Setting
IV-C Quantitative Analysis
Tables II-V show the performance compared to six state-of-the-art methods in three widely used datasets, which are regarded as three different domains. The model is trained in two selected domains and tested on the remaining one to evaluate cross-domain generalization. Consistent with established protocols, we express the ETH-UCY and NBA results in meters and the SDD results in pixels. To eliminate the impact of different units of measurement on the model, we unified them as meters during training. Given our emphasis on generalization ability and a distinct experimental setup compared to previous state-of-the-art methods, we reproduced previous state-of-the-art methods under the generalization setting through their officially released codes without post-processing to ensure fair comparisons.
ETH-UCY. Table II shows the generalization performance on the ETH-UCY dataset, which is trained on the SDD and NBA source domains. The experimental results show that our method achieves the best performance in both the ADE and FDE metrics. Specifically, our method improves the performance of the previous best method, Graph-TERN [63], from $0.36$ to $0.23$ on average ADE and from $0.60$ to $0.42$ on average FDE. Besides, we also outperform other approaches in all five subsets on both ADE and FDE metrics, which demonstrates our superior generalization capability. The underlying reason is that our recurrent alignment strategy enables the model to learn domain-invariant features across different scenarios.
SDD. Table III shows the generalization performance on the SDD dataset, which is trained on the ETH-UCY and NBA source domains. The experimental results show that our method achieves the best performance in both the ADE and FDE metrics compared to the other methods. Specifically, our method improves performance over the previous best method, Graph-TERN [63], from $14.76$ to $10.97$ in the average ADE metric and from $25.47$ to $19.95$ in average FDE metric. This demonstrates that our method has superior generalization capability compared to other approaches.
NBA. Table V shows the generalization performance on the NBA dataset, which is trained on ETH-UCY and SDD source domains. The experimental results show that our method achieves the best performance in both ADE and FDE metrics. Specifically, our method improves the average ADE/FDE from $1.28/2.64$ to $2.11/4.45$ compared to the previous best performance method, SocialVAE [1]. Notably, NBA dataset is captured in the basketball court, which presents denser interactions and a larger domain gap compared to the other datasets. Our method achieves more performance improvement facing such a larger domain gap, i.e. $44.8\%/45.3\%$ (NBA), $36.1\%/30\%$ (ETH-UCY) and $25.6\%/21.6\%$ (SDD), compared to Graph-TERN [63]. These results underscore our method’s exceptional ability to generalize across varied and complex domains for pedestrian trajectory prediction.